# The interplay of WGC and WCCC via charged scalar field fluxes in the RPST framework

### v1

@@ Abstract
In this paper, we investigate the weak cosmic censorship conjecture (WCCC) for the Reissner-Nordstrom (R-N) AdS black hole in a restricted phase space thermodynamics (RPST). Also here, we consider energy flux and equivalence mass-energy principle and examine the weak gravity conjecture (WGC) and the weak cosmic censorship conjecture. The incoming and outgoing energy flux leads to changes in the black hole. In that case, by applying the first law, we examined whether the second law of thermodynamics is valid. And also one can say that, in the case where absorption and superradiance are in the saturated to an equilibrium. Also, by using the thermodynamics of black holes in the restricted phase space, we show that if the black hole is in an extreme or close to an extreme state with radiation and particle absorption, the weak cosmic censorship conjecture is established. In addition, with the help of equivalence mass and energy principle and second-order approximation, in the near extremity, we find that when the black hole radiates and its central charge is greater than the scaled electric charge, the superradiance particles obey the weak gravity conjecture, and this causes the black hole to move further away from its extreme state. But when the particles that obey the weak gravity conjecture are attracted to the black hole when the black hole is very small. Then, in this case we note that the black hole becomes closer to its extreme state.
Keywords: R-N AdS Black Hole, Weak Gravity Conjecture, Weak Cosmic Censorship Conjecture, Restricted Phase Space.
@@ 1 Introduction
Many attempts are made to establish a connection between quantum theory and gravity. Among these efforts is the swampland program, used for the phenomenology of quantum gravity (string theory). In the swampland program, effective field theories that are compatible with quantum gravity are considered landscape and those that are not compatible are considered swampland [1, 2, 3]. Also, one of the important conjectures of this swampland program, which defines the border between the landscape and the swampland, is the weak gravity conjecture, which is considered the upper limit of mass, hence gravity is the weakest force [4, 5]. A lot of work has been done in this field [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. One of the most important phenomena in cosmology is the black hole. Since the black hole is the solution to Einstein’s equation, we can use the weak gravity conjecture derived from the string theory for the black hole to find a connection between these two theories.
@@
So, the above information give us motivation to arrange paper as: In section 2 we show in extended phase space, the cosmological constant is equivalence to thermodynamic pressure and black hole’s volume as its conjugate thermodynamic variable. Also, in this space, the mass of the black hole is regarded as enthalpy. Also in this section, we study the first law of thermodynamics and Smarr for the charged AdS black hole in the extended phase space. By using dual theory of holography, the central charge, also has a conjugate thermodynamic quantity as a chemical potential. We study the first law of thermodynamics and Smarr’s relation for a charged AdS black hole in restricted phase space. In section 3 , we generally examine the charged scalar field near the event horizon of corresponding black hole. In this study, we examine the scattering behavior of the coupled massive scalar field possessing an electric charge interacting with the charged AdS black hole within a 4-dimensional spacetime framework. The charged massive scalar field in charged Ads black hole is discussed in section 4. In section 5, we investigate WCCC in near-extremal and extremal charged AdS black hole with the above mentioned information. In section 6 we study the WGC for the corresponding black hole with help of equivalence of mass and energy principle. Finally we have some results and conclusion in section 7.
@@ 2 RPS thermodynamics and charged Ads black hole
In the extended phase space, the cosmological constant $(\Lambda)$ is equivalence to thermodynamic pressure of black hole. As we know, the change of AdS radius $(\ell)$ will change the black hole’s thermodynamic pressure, with the black hole’s volume as its conjugate thermodynamic variable [57],
$\begin{split}P=-\frac{\Lambda}{8\pi}=\frac{(d-1)(d-2)}{16\pi\ell^{2}},\qquad V=\bigg{(}\frac{\partial M}{\partial P}\bigg{)}_{S,Q},\end{split}$ (1)
@@
$\begin{split}M=E+PV,\end{split}$ (2)
@@
$\begin{split}dM=TdS+\Phi dQ+VdP\end{split}$ (3)
@@
$\begin{split}M=\frac{d-2}{d-3}TS+\Phi Q-\frac{2}{d-3}PV,\end{split}$ (4)
here, $S$, $\Phi$ and $T$ are the entropy, electric potential and Hawking temperature of the black hole respectively. In restricted phase space, Newton’s constant $(G)$ varies and the AdS radius $(\ell)$ is considered fixed, thus the black hole’s mass is seen as its internal energy $(E)$ [44, 45, 49, 54],
$\begin{split}E=M.\end{split}$ (5)
By using dual theory of holography, the central charge $(C)$ as an extensive variable, also it has conjugate thermodynamic quantity as a chemical potential $(\mu)$ [44, 45, 43],
$\begin{split}C=\frac{\ell^{d-2}}{G},\qquad\mu=\bigg{(}\frac{\partial E}{\partial C}\bigg{)}_{S,\hat{Q}},\end{split}$ (6)
@@
$\begin{split}\hat{Q}=Q\sqrt{C},\qquad\hat{\Phi}=\frac{\Phi(r_{h})}{\sqrt{C}}=\frac{\hat{Q}}{r_{h}C}\end{split}$ (7)
@@
$\begin{split}dE=TdS+\hat{\Phi}d\hat{Q}+\mu dC\end{split}$ (8)
@@
$\begin{split}E=TS+\hat{\Phi}\hat{Q}+\mu C\end{split}$ (9)
Next, we obtain the metric, entropy, mass, and temperature of the charged black hole in the restricted phase space. So, in that case, the metric background of charged AdS black hole in 4-dimension will be following form,
$\begin{split}ds^{2}=-h(r)dt^{2}+\frac{1}{h(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}\end{split}$ (10)
@@
$\begin{split}h(r)=1-\frac{2GM}{r}+\frac{GQ^{2}}{r^{2}}+\frac{r^{2}}{\ell^{2}}\qquad or\qquad h(r)=1-\frac{2\ell^{2}M}{Cr}+\frac{\ell^{2}\hat{Q}^{2}}{C^{2}r^{2}}+\frac{r^{2}}{\ell^{2}}\end{split}$ (11)
here, we assume units as $k_{B}=c=\hbar=1$. When $\frac{Q}{M}<\sqrt{G}$, the black hole has two internal and external event horizons, we denote the external event horizon by $r_{h}$. Also, in the case of $\frac{Q}{M}=\sqrt{G}$, the black hole is in an extreme state with only one event horizon. Also, the entropy of charged black hole in the corresponding space is obtained by following equation,
$\begin{split}S=\frac{\pi r_{h}^{2}}{G},\qquad r_{h}=\ell\bigg{(}\frac{S}{\pi C}\bigg{)}^{\frac{1}{2}}\end{split}$ (12)
By using equations (7),(11) and (12), one can obtain the mass of black hole as,
$M=\frac{\pi CS+\pi^{2}\hat{Q}^{2}+S^{2}}{2\pi^{3/2}\ell\sqrt{CS}}$ (13)
@@
$T=\bigg{(}\frac{\partial M}{\partial S}\bigg{)}_{\hat{Q},C}=\frac{\pi CS-\pi^{2}\hat{Q}^{2}+3S^{2}}{4\pi^{3/2}\ell S\sqrt{CS}}$ (14)
@@
$\mu=\bigg{(}\frac{\partial M}{\partial C}\bigg{)}_{S,\hat{Q}}=\frac{\pi CS-\pi^{2}\hat{Q}^{2}-S^{2}}{4\pi^{3/2}C\ell\sqrt{CS}}$ (15)
@@ 3 Charged massive scalar field in charged Ads black hole
Black holes are capable of acquiring conserved quantities, including energy, momenta and electric charge, through the interaction with an external field. The obtained conserved quantities from black hole corresponds to the fluxes of the external field. Through these fluxes, the black hole undergoes alterations in its state as it engages with the external field. By analyzing the fluxes, we can approximate the transformations occurring within the black hole over infinitesimal time intervals. In this study, we examine the scattering behavior of the coupled massive scalar field possessing an electric charge interacting with the charged AdS black hole within a 4-dimensional spacetime framework. The equation of the massive charged scalar field near the charged black hole is given by [36, 58, 59],
$\frac{1}{\sqrt{-g}}\partial_{\mu}(g^{\mu\nu}\sqrt{-g}\partial_{\nu}\Psi)-2iqg^{\mu\nu}A_{\mu}\partial_{\nu}\Psi-q^{2}g^{\mu\nu}A_{\mu}A_{\nu}\Psi-\mu_{s}^{2}\Psi=0$ (16)
@@
$\Psi(t,r,\theta,\phi)=e^{-i\omega t}R(r)Y_{lm}(\theta,\phi)$ (17)
where $R(r)$, $Y_{lm}(\theta,\phi)$ and $\omega$ are radial functions, spherical harmonic functions and energy of particle, respectively. Furthermore, the integer parameters $l$ and $m$ denote the spherical and azimuthal harmonic indices of the resonant eigenvalues that define the characteristics of the charged massive scalar fields within the charged black-hole spacetime. We get two radial and angular parts equations by putting relations (10) and (17) into equation (16),
$h(r)\frac{d}{dr}\bigg{[}r^{2}h(r)\frac{dR(r)}{dr}\bigg{]}+\bigg{[}r^{2}\bigg{(}\omega-\frac{qQ}{r}\bigg{)}-h(r)\bigg{(}l(l+1)+\mu_{s}r^{2}\bigg{)}\bigg{]}R(r)=0$ (18)
$\bigg{[}\frac{1}{\sin^{2}\theta}\frac{\partial^{2}}{\partial\phi^{2}}+\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\bigg{(}\sin\theta\frac{\partial}{\partial\theta}\bigg{)}\bigg{]}Y_{lm}(\theta,\phi)=-l(l+1)Y_{lm}(\theta,\phi)$ (19)
@@
$\frac{d^{2}R}{dy^{2}}+\frac{2h(r)}{r}\frac{dR}{dy}+\bigg{[}(\omega-\frac{qQ}{r})^{2}-\frac{h(r)}{r^{2}}\bigg{(}l(l+1)-\mu_{s}r^{2}\bigg{)}\bigg{]}R=0$ (20)
Since the flux of the scalar field into the black hole is important, we check equation (20) near the event horizon of the black hole $(r\rightarrow r_{h})$, in which case we have a Schrödinger-like equation,
$\frac{d^{2}R}{dy^{2}}+\bigg{(}\omega-\frac{qQ}{r_{h}}\bigg{)}^{2}R=0$ (21)
@@
$R(r)=e^{\pm i(\omega-\frac{qQ}{r_{h}})^{2}y}=e^{\pm i[\omega-q\Phi(r_{h})]^{2}y}$ (22)
Here, +/+/-+ / - sign indicates the outgoing and ingoing of the fields. Therefore, the wave function of the charged scalar field near the event horizon of the charged black hole takes the following form,
$\Psi(t,r,\theta,\phi)=e^{-i\omega t}e^{\pm i[\omega-q\Phi(r_{h})]^{2}y}Y_{lm}(\theta,\phi)$ (23)
@@
$T^{\mu}_{\nu}=\frac{1}{2}\mathcal{D}^{\mu}\Psi\partial_{\nu}\Psi^{*}+\frac{1}{2}\mathcal{D}^{*\mu}\Psi^{*}\partial_{\nu}\Psi-\delta^{\mu}_{\nu}\bigg{(}\frac{1}{2}\mathcal{D}_{\rho}\Psi\mathcal{D}^{*\rho}\Psi^{*}-\frac{\mu_{s}^{2}}{2}\Psi\Psi^{*}\bigg{)}$ (24)
@@
$j^{\alpha}=\frac{iq}{2}\Psi\mathcal{D}^{*\alpha}\Psi^{*}-\frac{iq}{2}\Psi^{*}\mathcal{D}^{\alpha}\Psi$ (25)
Therefore, from equations (24) and (25), the energy flux and electric charge near the event horizon are obtained by [36],
@@
From the above equations, when $\omega>q\Phi_{h}$, the energy and charge flux is positive, indicating the particle’s entry into the black hole. But for $\omega<q\Phi_{h}$, the energy and charge flux become negative, which means the exit of the particle from the black hole, and this phenomenon is called black hole superradiation. These fluxes can change the internal energy and electric charge of black holes in infinitesimal time interval $dt$. Since we consider the thermodynamics of the black hole in the RPS, the changes in the internal energy of the black hole are equal to the changes in the mass of the black hole. Also, we consider the changes in electric charge according to equations (7) and (27), which has been re-scaled in this space. So in that case, we have,
$dE=dM=\omega(\omega-\hat{q}\hat{\Phi}_{h})r_{h}^{2}dt$ (28)
$d\hat{Q}=\hat{q}(\omega-\hat{q}\hat{\Phi}_{h})r_{h}^{2}dt+\frac{\hat{Q}}{2C}dC$ (29)
@@
$\hat{q}=q\sqrt{C},\qquad\hat{\Phi}_{h}=\frac{\Phi_{h}}{\sqrt{C}}$ (30)

### v2

@@ Abstract
In this paper, we investigate the weak cosmic censorship conjecture (WCCC) for the Reissner-Nordström (R-N) AdS black hole within the framework of restricted phase space thermodynamics (RPST). Additionally, we consider energy flux and the equivalence mass-energy principle to examine the weak gravity conjecture (WGC) alongside the weak cosmic censorship conjecture. The interaction of incoming and outgoing energy flux leads to changes in the black hole. We examine whether the second law of thermodynamics holds in this scenario. It is noted that, when absorption and superradiance reach equilibrium, and by using the thermodynamics of black holes in restricted phase space, if the black hole is in or near an extreme state with radiation and particle absorption, the weak cosmic censorship conjecture is upheld. Furthermore, by employing the equivalence mass-energy principle and second-order approximation near extremality, we find that when the black hole radiates and its central charge exceeds the scaled electric charge, the superradiance particles adhere to the weak gravity conjecture. This causes the black hole to move away from its extreme state. However, when particles that obey the weak gravity conjecture are attracted to a very small black hole, the black hole approaches its extremality state.
Keywords: Reissner-Nordström AdS Black Hole, Weak Gravity Conjecture, Weak Cosmic Censorship Conjecture, Restricted Phase Space
@@ 1 Introduction
Many attempts are made to establish a connection between quantum theory and gravity. Among these efforts is the swampland program, used for the phenomenology of quantum gravity (string theory). In the swampland program, effective field theories compatible with quantum gravity are considered landscape and those not compatible are considered swampland [1, 2, 3]. Also, one of the important conjectures of this swampland program, which defines the border between the landscape and the swampland, is the weak gravity conjecture, which is considered the upper limit of mass, hence gravity is the weakest force [4, 5]. Also, different cosmological concepts such as inflation, dark energy, black holes physics, thermodynamics, etc investigated with the various conjectures of the swampland program. For further study, you can see in [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. One of the most important phenomena in cosmology is the black hole. Since the black hole is the solution to Einstein’s equation, we can use the weak gravity conjecture derived from the string theory for the black hole to find a connection between these two theories.
@@
So, the above information gives us motivation to arrange the paper as: In section 2 we show in extended phase space, the cosmological constant is equivalence to thermodynamic pressure and black hole’s volume as its conjugate thermodynamic variable. Also, in this space, the mass of the black hole is regarded as enthalpy. Also in this section, we study the first law of thermodynamics and Smarr for the charged AdS black hole in the extended phase space. By using the dual theory of holography, the central charge, also has a conjugate thermodynamic quantity as a chemical potential. We study the first law of thermodynamics and Smarr’s relation for a charged AdS black hole in restricted phase space. In section 3, we generally examine the charged scalar field near the event horizon of the corresponding black hole. In this study, we examine the scattering behavior of the coupled massive scalar field possessing an electric charge interacting with the charged AdS black hole within a 4-dimensional spacetime framework. The charged massive scalar field in charged Ads black hole is discussed in section 4. In section 5, we investigate WCCC in near-extremal and extremal-charged AdS black holes with the above-mentioned information. In section 6 we study the WGC for the corresponding black hole with the help of the equivalence of mass and energy principle. Finally, we explain the conclusions in section 7.
@@ 2 RPS thermodynamics and charged AdS black hole
In the extended phase space, the cosmological constant $(\Lambda)$ is equivalence to the thermodynamic pressure of a black hole. As we know, the change of AdS radius $(\ell)$ will change the black hole’s thermodynamic pressure, with the black hole’s volume as its conjugate thermodynamic variable [97],
$\begin{split}P=-\frac{\Lambda}{8\pi}=\frac{(d-1)(d-2)}{16\pi\ell^{2}},\qquad V=\bigg{(}\frac{\partial M}{\partial P}\bigg{)}_{S,Q},\end{split}$ (2.1)
@@
$\begin{split}M=E+PV,\end{split}$ (2.2)
@@
$\begin{split}dM=TdS+\Phi dQ+VdP\end{split}$ (2.3)
@@
$\begin{split}M=\frac{d-2}{d-3}TS+\Phi Q-\frac{2}{d-3}PV,\end{split}$ (2.4)
here, $S$, $\Phi$, and $T$ are the entropy, electric potential, and Hawking temperature of the black hole respectively. In restricted phase space, Newton’s constant $(G)$ varies and the AdS radius $(\ell)$ is considered fixed, thus the black hole’s mass is seen as its internal energy $(E)$ [84, 85, 89, 94],
$\begin{split}E=M.\end{split}$ (2.5)
By using the dual theory of holography, the central charge $(C)$ as an extensive variable has conjugate thermodynamic quantity as a chemical potential $(\mu)$ [84, 85, 83],
$\begin{split}C=\frac{\ell^{d-2}}{G},\qquad\mu=\bigg{(}\frac{\partial E}{\partial C}\bigg{)}_{S,\hat{Q}},\end{split}$ (2.6)
@@
$\begin{split}\hat{Q}=Q\sqrt{C},\qquad\hat{\Phi}=\frac{\Phi(r_{h})}{\sqrt{C}}=\frac{\hat{Q}}{r_{h}C}\end{split}$ (2.7)
@@
$\begin{split}dE=TdS+\hat{\Phi}d\hat{Q}+\mu dC\end{split}$ (2.8)
@@
$\begin{split}E=TS+\hat{\Phi}\hat{Q}+\mu C\end{split}$ (2.9)
Next, we obtain the metric, entropy, mass, and temperature of the charged black hole in the restricted phase space. So, in that case, the metric background of the charged AdS black hole in 4-dimension will be the following form,
$\begin{split}ds^{2}=-h(r)dt^{2}+\frac{1}{h(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}\end{split}$ (2.10)
@@
$\begin{split}h(r)=1-\frac{2GM}{r}+\frac{GQ^{2}}{r^{2}}+\frac{r^{2}}{\ell^{2}}\qquad or\qquad h(r)=1-\frac{2\ell^{2}M}{Cr}+\frac{\ell^{2}\hat{Q}^{2}}{C^{2}r^{2}}+\frac{r^{2}}{\ell^{2}}.\end{split}$ (2.11)
Here, we assume units as $k_{B}=c=\hbar=1$. When $\frac{Q}{M}<\sqrt{G}$, the black hole has two internal and external event horizons, we denote the external event horizon by $r_{h}$. Also, in the case of $\frac{Q}{M}=\sqrt{G}$, the black hole is in an extreme state with only one event horizon. Also, the entropy of the charged black hole in the corresponding space is obtained by the following equation,
$\begin{split}S=\frac{\pi r_{h}^{2}}{G},\qquad r_{h}=\ell\bigg{(}\frac{S}{\pi C}\bigg{)}^{\frac{1}{2}}\end{split}$ (2.12)
By using equations (2.7),(2.11) and (2.12), one can obtain the mass of black hole as,
$M=\frac{\pi CS+\pi^{2}\hat{Q}^{2}+S^{2}}{2\pi^{3/2}\ell\sqrt{CS}}$ (2.13)
@@
$T=\bigg{(}\frac{\partial M}{\partial S}\bigg{)}_{\hat{Q},C}=\frac{\pi CS-\pi^{2}\hat{Q}^{2}+3S^{2}}{4\pi^{3/2}\ell S\sqrt{CS}}$ (2.14)
@@
$\mu=\bigg{(}\frac{\partial M}{\partial C}\bigg{)}_{S,\hat{Q}}=\frac{\pi CS-\pi^{2}\hat{Q}^{2}-S^{2}}{4\pi^{3/2}C\ell\sqrt{CS}}$ (2.15)
@@ 3 Charged massive scalar field in charged AdS black hole
Black holes are capable of acquiring conserved quantities, including energy, momenta, and electric charge, through the interaction with an external field. The obtained conserved quantities from the black hole correspond to the fluxes of the external field. Through these fluxes, the black hole undergoes alterations in its state as it engages with the external field. By analyzing the fluxes, we can approximate the transformations occurring within the black hole over infinitesimal time intervals. In this study, we examine the scattering behavior of the coupled massive scalar field possessing an electric charge interacting with the charged AdS black hole within a 4-dimensional spacetime framework. The equation of the massive charged scalar field near the charged black hole is given by [76, 98, 99],
$\frac{1}{\sqrt{-g}}\partial_{\mu}(g^{\mu\nu}\sqrt{-g}\partial_{\nu}\Psi)-2iqg^{\mu\nu}A_{\mu}\partial_{\nu}\Psi-q^{2}g^{\mu\nu}A_{\mu}A_{\nu}\Psi-\mu_{s}^{2}\Psi=0$ (3.1)
@@
$\Psi(t,r,\theta,\phi)=e^{-i\omega t}R(r)Y_{lm}(\theta,\phi)$ (3.2)
where $R(r)$, $Y_{lm}(\theta,\phi)$ and $\omega$ are radial functions, spherical harmonic functions, and energy of the particle, respectively. Furthermore, the integer parameters $l$ and $m$ denote the spherical and azimuthal harmonic indices of the resonant eigenvalues that define the characteristics of the charged massive scalar fields within the charged black-hole spacetime. We get two radial and angular parts equations by putting relations (2.10) and (3.2) into equation (3.1),
$h(r)\frac{d}{dr}\bigg{[}r^{2}h(r)\frac{dR(r)}{dr}\bigg{]}+\bigg{[}r^{2}\bigg{(}\omega-\frac{qQ}{r}\bigg{)}-h(r)\bigg{(}l(l+1)+\mu_{s}r^{2}\bigg{)}\bigg{]}R(r)=0$ (3.3)
$\bigg{[}\frac{1}{\sin^{2}\theta}\frac{\partial^{2}}{\partial\phi^{2}}+\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\bigg{(}\sin\theta\frac{\partial}{\partial\theta}\bigg{)}\bigg{]}Y_{lm}(\theta,\phi)=-l(l+1)Y_{lm}(\theta,\phi)$ (3.4)
@@
$\frac{d^{2}R}{dy^{2}}+\frac{2h(r)}{r}\frac{dR}{dy}+\bigg{[}(\omega-\frac{qQ}{r})^{2}-\frac{h(r)}{r^{2}}\bigg{(}l(l+1)-\mu_{s}r^{2}\bigg{)}\bigg{]}R=0$ (3.5)
Since the flux of the scalar field into the black hole is important, we check equation (3.5) near the event horizon of the black hole $(r\rightarrow r_{h})$, in which case we have a Schrödinger-like equation,
$\frac{d^{2}R}{dy^{2}}+\bigg{(}\omega-\frac{qQ}{r_{h}}\bigg{)}^{2}R=0$ (3.6)
@@
$R(r)=e^{\pm i(\omega-\frac{qQ}{r_{h}})^{2}y}=e^{\pm i[\omega-q\Phi(r_{h})]^{2}y}$ (3.7)
Here, the +/+/-+ / - sign indicates the outgoing and ingoing of the fields. Therefore, the wave function of the charged scalar field near the event horizon of the charged black hole takes the following form,
$\Psi(t,r,\theta,\phi)=e^{-i\omega t}e^{\pm i[\omega-q\Phi(r_{h})]^{2}y}Y_{lm}(\theta,\phi)$ (3.8)
@@
$T^{\mu}_{\nu}=\frac{1}{2}\mathcal{D}^{\mu}\Psi\partial_{\nu}\Psi^{*}+\frac{1}{2}\mathcal{D}^{*\mu}\Psi^{*}\partial_{\nu}\Psi-\delta^{\mu}_{\nu}\bigg{(}\frac{1}{2}\mathcal{D}_{\rho}\Psi\mathcal{D}^{*\rho}\Psi^{*}-\frac{\mu_{s}^{2}}{2}\Psi\Psi^{*}\bigg{)}$ (3.9)
@@
$j^{\alpha}=\frac{iq}{2}\Psi\mathcal{D}^{*\alpha}\Psi^{*}-\frac{iq}{2}\Psi^{*}\mathcal{D}^{\alpha}\Psi$ (3.10)
Therefore, from equations (3.9) and (3.10), the energy flux and electric charge near the event horizon are obtained by [76],
@@
From the above equations, when $\omega>q\Phi_{h}$, the energy and charge flux are positive, indicating the particle’s entry into the black hole. But for $\omega<q\Phi_{h}$, the energy and charge flux becomes negative, which means the exit of the particle from the black hole, and this phenomenon is called black hole superradiation. These fluxes can change the internal energy and electric charge of black holes in infinitesimal time interval $dt$. Since we consider the thermodynamics of the black hole in the RPS, the changes in the internal energy of the black hole are equal to the changes in the mass of the black hole. Also, we consider the changes in electric charge according to equations (2.7) and (3.12), which have been re-scaled in this space. So in that case, we have,
$dE=dM=\omega(\omega-\hat{q}\hat{\Phi}_{h})r_{h}^{2}dt$ (3.13)
$d\hat{Q}=\hat{q}(\omega-\hat{q}\hat{\Phi}_{h})r_{h}^{2}dt+\frac{\hat{Q}}{2C}dC$ (3.14)
@@
$\hat{q}=q\sqrt{C},\qquad\hat{\Phi}_{h}=\frac{\Phi_{h}}{\sqrt{C}}$ (3.15)